triangle ABC is a right triangle.If AC = 5 and side BC = 6, what is the measure of side AB?
Answer by nyc_function(2741)

Use the Pythagorean Theorem: a^2 + b^2 = c^2

(AC)^2 + (BC)^2 = (AB)^2
5^2 + 6^2 = (AB)^2
25 + 36 = (AB)^2
61 = (AB)^2
sqrt{61} = AB
By the way, sqrt = square root for short.